Find two consecutive integers such that the sum of the smaller and twice the greater is 32
Answer by Maths68(1474)

Let
Smaller integers = x
Greater = x+1
==========================================
Given
x+2(x+1)=32
x+2x+2=32
3x=32-2
3x=30
x=30/3
x=10
================================================
Smaller integer = x = 10
Greater integer = x+1 = 10+1=11
=================================================
Check
x+2(x+1)=32
10+2(10+1)=32
10+2(11)=32
10+22=32
32=32